A 5 kg plastic tank that has a volume of 0.2 m3 is filled with liquid water. Assuming the density of water is 1000 kg/m3, determine the mass in kg of the combined system


1
Expert's answer
2021-08-17T07:00:02-0400

mass of water = density × volume

mass of water = 1000 × 0.2 = 200kg


mass of tank = 5kg


Total mass = mass of water + mass of tank = "200 + 5" = "205\ kg"